OpenStudy (anonymous):

The displacement of a particle on a vibrating string is given by the equation s(t)=11+1/4sin(14PIt), where s is measured in centimeters and in seconds. Find the velocity of the particle after t seconds.

OpenStudy (anonymous):

Find the first derivative of the given expression.

OpenStudy (anonymous):

am unsure how to do that, step by step help would be great

OpenStudy (anonymous):

I sit the expression? \[s(t)=11+\frac{ 1 }{ 4\sin(14{\pi}t) }\]

OpenStudy (anonymous):

Is it the expression?

OpenStudy (anonymous):

s(t)= (1/4) sin(14pit)

OpenStudy (anonymous):

You mean this ?????\[s(t)=\frac{ 1 }{ 4 }\sin(14{\pi}t)\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[s^{'}(t)=v(t)=\frac{ 7 }{ 2 }{\pi}\cos(14{\pi}t)\]

OpenStudy (anonymous):

hell yeah...thanks!